OpenStudy (anonymous):

how to solve this exponential equation? 5(x-6)=25(x-6)

OpenStudy (mertsj):

I don't see any exponents

OpenStudy (mertsj):

\[5^{x-6}=25^{x-6}\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yeaaaah!

OpenStudy (anonymous):

It's x-4 though.

OpenStudy (mertsj):

\[5^{x-5}=5^{2x-12}\]

OpenStudy (mertsj):

So x-6=2x-12 and x = 6

OpenStudy (anonymous):

Ah? \[5^{x-4}=25^{x-6}\]

OpenStudy (anonymous):

?